OpenStudy (anonymous):

Let u = <7, -3>, v = <-9, 5>. Find 4u - 3v.

OpenStudy (anonymous):

4<7,-3> = <28,-12> 3<-9,5> = <-27, 15> Then subtract 28-(-27)=55 -12-15=-27 <55,-27> You are just multiplying by the magnitude of the vector

OpenStudy (anonymous):

thank you!